# Advection of the image point in probabilistically-reconstructed phase spaces

### v2

@@ 3.4 NEMO model
Figure 14: The same as in Fig. 12 but for SRV. Note that adding more EOFs makes the time mean energy spectral density, $\langle ESD\rangle$, closer to the reference one; compare the $\langle ESD\rangle$ of the HP solution for $m=200$ (red) and $m=400$ (magenta) with the reference (black). The $\langle ESD\rangle$ of the modelled solution (blue) is much lower compared with that of the HP solution, especially for smaller scales.
4 Conclusions
@@
In this study we have proposed to use the joint probability distribution (JPD) to probabilistically reconstruct the phase space which can be used when the original reference phase space (computed from numerical solutions, observations, or both) is too sparse to allow the use of data-driven methods. The probabilistically-reconstructed phase space is calculated by sampling from the JPD, thus providing an extra source of data from the reference distribution and ensuring that both the reference phase space and the reconstructed one share the same reference distribution. We have shown how the probabilistic reconstruction works on the example of Chua’s system and then successfully applied it within the context of the global ocean model NEMO.
@@
The current methodology is limited to the phase space of HP solution computed in the probabilistically-reconstructed phase space. The phase space of HP solution is larger than the phase space computed from the data distribution (i.e. the probabilistically-reconstructed one), because the latter might have voids (which cannot be patched up by sampling from the JPD), while the former has no voids. Giving these limitations, the proposed methodology does allow us to compute the HP solution in the probabilistically-reconstructed (and reduced) phase space which is more accurate compared with the 1/4-NEMO simulation nonetheless. However, if the reference distribution changes over time (e.g. the reference solution drifts away from the phase space, used to compute the reference distribution) then the HP solution cannot follow these changes.
@@
We have also found that the HP method can be used directly in reduced phase spaces thus allowing several-orders-of-magnitude acceleration (compared to the 1/4-resolution NEMO run), which can be translated into using larger ensemble of HP solutions for probabilistic predictions. Note that we compare only the degrees of freedom used in the HP method and those used in the NEMO model to get the solution at $1/4^{\circ}$ resolution. It goes without saying that the NEMO model evolves much more variables than the HP method, but in this comparison we pretend that NEMO needs only $\sim$30000 grid points to compute the solution, while the HP method uses only 200 in the compressed phase space, and more importantly gives a solution which is closer to the reference one than that of the $1/4^{\circ}$ NEMO simulation.
@@
We have studied how the proposed method performs on corrupted reference data (for which we considered three cases: randomly-missed states, gaps, and cuts) and found that the method works well even when sampling from the JPD cannot patch up voids in the reference phase space. It allows us to conclude that the HP method in probabilistically-reconstructed phase spaces can be used as a fast reanalysis in which the complex dynamics of a comprehensive ocean model is replaced with the HP solution, as well as a dynamic interpolation method to stitch gaps in observational data.
@@
For studying climate change, one might think that the reference data set might be a limiting factor. And indeed, it is, but only if future climate states are not within the reference phase space. As it is a priori unknown whether future climate states are located in the reference phase space, one might consider using this methodology as an alternative to parameterisations in comprehensive models for near-term climate predictions (NTCP), instead of using it as a tool for forecasting, since more research is needed to address this question. Therefore, the next step in this direction would be to assess prediction skills and achievable forecast range of the method alone and in combination with high-resolution forecasting systems for NTCP.
@@
The output of the global 1/4 (ORCA025-N4001) and 1/12 (ORCA0083-N006) resolution NEMO model used in this study is available on JASMIN ( https://jasmin.ac.uk/users/access/ ) from these locations:
/gws/nopw/j04/nemo_vol1/ORCA025-N401 and
/gws/nopw/j04/nemo_vol1/ORCA0083-N006 .
@@ Acknowledgments
The author thanks the Natural Environment Research Council for the support of this work through the projects CLASS and ATLANTIS (P11742-01), as well as Andrew Coward and Chris Wilson for the production of and help with the NEMO datasets, respectively.

### v3

@@ 3.4 NEMO model
Figure 14: The same as in figure 12 but for SRV; the units in subplots (a)-(d) and (e) are ${\rm s}^{-1}]$ and ${\rm s}^{-1}{\rm deg}^{2}]^{2}$, respectively. Note that adding more EOFs makes the time mean energy spectral density, $\langle ESD\rangle$, closer to the reference one; compare the $\langle ESD\rangle$ of the HP solution for $m=200$ (red) and $m=400$ (magenta) with the reference (black). The $\langle ESD\rangle$ of the modelled solution (blue) is much lower compared with that of the HP solution, especially for smaller scales.
3.5 Damaged reference data
Damaged reference data sets are frequently encountered in observations used for data assimilation and reanalysis in comprehensive ocean models. There is an extensive toolkit of methods currently in circulation to repair damaged data so that it could be of further use in computations. As we have seen in the example of Chua’s circuit, the HP method in probabilistically-reconstructed phase spaces can work directly with unreliable reference data sets thus requiring no extra tools to cope with this issue. In this section, we study how damaged reference data from the NEMO model impacts the HP method. Similar to the scenario with Chua’s circuit, we examine the reference phase space damaged by randomly missed states (D1), gaps (D2), and cuts (D3). Each damaged data set is of the size of the validation period to ensure its length is the same as the one used in runs with undamaged data above. We only focus on the surface relative vorticity, as it is more sensitive to errors than temperature. The results show that regardless of the nature of damage, the proposed methodology outperforms the modelled solution by providing a more accurate representation of the time-mean flow (figure 15), the root mean square error of the HP solution is comparable to that of the modelled solution. Furthermore, the energy spectrum of the HP solution (across all types of damage) is closer to the reference spectrum compared with the one of the modelled solution (figure 16). These findings allow us to conclude that the proposed methodology can be used as a dynamic interpolation method for repairing damaged observational data.
The HP method, in its standard form, is not designed for within-field spatial gap-filling, which however can be present in real observations (for example, some regions can be masked by clouds). It assumes complete phase-space inputs at each time step and operates on the temporal evolution of full system states, rather than performing grid-level inpainting in physical space. While the HP method can, in principle, be adapted to handle partial fields (e.g., by operating on local patches or integrating with a statistical imputation step), such extensions are beyond the scope of the present study. Here we focus on demonstrating the HP method for working in phase spaces with damaged structure.
Figure 15: Shown are (a) difference between the time mean reference SRV and modelled SRV, as well as HP SRV (D1), HP SRV (D2), HP SRV (D3) (top to bottom); (b) root mean square error between the reference SRV and modelled SRV, as well as HP SRV (D1), HP SRV (D2), HP SRV (D3) (top to bottom). We use m=200 in the EOF-PC decomposition for (a) and (b). The time mean SRV of the HP solution computed from the damaged data ((a) – second to fourth row) is more accurate than that of the modelled solution ((a) – top), while the root mean square error of the HP solution is on par with that of the modelled solution (b).
Figure 16: Shown is the natural logarithm of the time mean energy spectral density $\langle ESD\rangle$ (in units of ${\rm s}^{-1}\cdot{\rm deg}^{2}]^{2}$) vs the magnitude of the wave vector $\mathbf{k}$. The $\langle ESD\rangle$ of the HP solution (red) and HP solutions computed from the damaged reference data sets (D1, D2, D3) are closer to the reference solution (black) than that of the modelled solution (blue). As in figure 15 we use m=200 in the EOF-PC decomposition.
4 Conclusions and Discussion
@@
In this study we have proposed to use the joint probability distribution (JPD) to probabilistically reconstruct the phase space when the original reference phase space (computed from numerical solutions, observations, or both) is too sparse for the HP approach to be effective. The probabilistically-reconstructed phase space is calculated by sampling from the JPD, which provides an additional source of data from the reference distribution. The method ensures that the reference phase space and the reconstructed space share the same reference distribution. We have shown how the probabilistic reconstruction works on the example of Chua’s system and then successfully applied it regionally within the context of the global ocean model NEMO.
On the scope and limitations of HP. The proposed methodology operates within the phase space defined by the HP solution computed in the probabilistically-reconstructed reference phase space. It is important to distinguish between these two objects. The probabilistically-reconstructed phase space is obtained by sampling from the reference data distribution and may still contain voids, particularly in regions where the original dataset lacked global support (such voids cannot be fully recovered by JPD sampling if entire regions of the attractor were never observed).
In contrast, the phase space traced by the HP solution is dynamically continuous: since the HP trajectory is obtained by integrating equations (1) or (4), it does not consist of isolated sampled points and therefore does not exhibit discrete voids. In this sense, the HP phase space may appear “larger” or more connected than the probabilistically-reconstructed cloud. However, its evolution remains intrinsically tied to the reference phase space, because at every time step the evolution of the image point relies on neighbourhood statistics computed from the reference dataset. The method therefore does not construct a global parametric model of the dynamics. Instead, it reconstructs trajectories using the reference phase space. Its generalization properties are therefore determined by the coverage of that phase space.
  • 1.
    Generalization to different initial conditions. If the initial condition lies within (or sufficiently close to) the reference phase space, the HP method can generate dynamically consistent trajectories. In this sense, the emulator generalizes across initial conditions belonging to the same dynamical regime represented in the reference data.
  • 2.
    Generalization to different physical parameters. The method does not extrapolate well beyond the parameter regime represented in the reference dataset. If the underlying physical parameters change in such a way that the attractor geometry undergoes significant structural changes, the reference phase space must be recomputed or augmented accordingly.
  • 3.
    Out-of-distribution states. If the trajectory enters regions far away from the reference phase space, the HP method cannot reliably reconstruct the dynamics, as the neighbourhood information becomes unrepresentative.
In this sense, the HP emulator should be viewed as a nonparametric, reference-phase-space-conditioned dynamical reconstruction tool rather than a globally trained parametric surrogate. Its predictive skill depends on phase-space coverage rather than parametric training.
@@
Despite these limitations, the proposed methodology enables the computation of the HP solution in the probabilistically-reconstructed (and also reduced) phase space, which proves to be more accurate than the 1/4-NEMO simulation. However, if the reference distribution changes over time (e.g. the reference solution drifts away from the phase space, used to compute the reference distribution) then the HP solution cannot follow these changes.
@@
We have also found that the HP method can be used directly in reduced phase spaces, thus leading to several-orders-of-magnitude acceleration (compared to the 1/4-resolution NEMO run). This acceleration can be translated into using larger ensembles of HP solutions for probabilistic predictions. Note that we compare only the degrees of freedom utilized in the HP method and in the NEMO model to get the solution at $1/4^{\circ}$ resolution. It goes without saying that the NEMO model evolves much more variables than the HP method, but in this comparison we pretend that NEMO needs only $\sim$30000 grid points to compute the solution, while the HP method uses only 200 in the compressed phase space, and more importantly gives a solution which is closer to the reference one than that of the $1/4^{\circ}$ NEMO simulation.
The computational cost of the proposed methodology is dominated by the offline preprocessing stage (PDF estimation and EOF decomposition). The HP trajectory integration in the reduced space is computationally inexpensive, involving only nearest-neighbour searches and local averaging.
@@
We have studied how the proposed method performs on damaged reference data (for which we considered three cases:randomly-missed states, gaps, and cuts) and found that the method works well even when sampling from the JPD cannot patch up voids in the reference phase space. It allows us to conclude that the HP method in probabilistically-reconstructed phase spaces can be used as a fast reanalysis method in which the complex dynamics of a comprehensive ocean model is replaced with the HP solution, as well as a dynamic interpolation method to fill gaps in observational data.
@@
When modelling climate changes with data-driven methods, one might assume that the reference data set could be a limiting factor. While it is true, it only becomes a concern if future climate states fall outside the reference phase space. Since it is a priori unknown whether future climate states are located in the reference phase space, it may be advisable to use this methodology as an alternative to parameterisations in comprehensive models for near-term climate predictions (NTCP), instead of using it solely as a forecasting tool. As with ensemble forecasts generated by physics-based models, an ensemble of HP-based emulations, each initialized from slightly different image points in phase space, will yield distinct trajectories and thus define a probabilistic description of the flow dynamics. However, the interpretation of this ensemble differs fundamentally from that of ensembles produced by physics-based models. In traditional physics-based models, ensemble members divergence arises from nonlinear instabilities, chaotic sensitivity to initial conditions, model error, variability in external forcing, etc. – all governed by dynamical equations. In contrast, HP ensembles reflect geometric sensitivity to the local structure of the reference phase space. Variability between HP members originates from how slightly perturbed initial conditions explore different regions of the phase space. Consequently, the HP ensemble spread reflects data-induced uncertainty, not dynamical chaos or forcing-driven divergence.
Although the HP method is formally time-agnostic and operates entirely within the geometry of a reference phase space, it can, under certain conditions, act as a forecast system. Specifically, if future states of the system lie within the statistical and geometric support of the reference phase space (i.e. they are already geometrically encoded into the reference phase space) then HP can evolve towards them. In this case, the ensemble reflects statistically admissible futures. As long as the evolution remains within the reference phase space, HP ensembles will closely resemble physics-based model ensembles in terms of time-evolving statistics such as mean, spread, spatial and temporal patterns.
However, while HP ensembles are computationally inexpensive and statistically consistent with sampled dynamics, they offer a geometric interpolation of uncertainty, rather than a physics-driven future forecast. HP cannot extrapolate beyond that; it lacks any mechanism to explore regions of phase space not already explicitly or implicitly represented in the reference data. This limitation is not unique to HP. It applies broadly to data-driven emulators that lack explicit knowledge of governing equations or external forcings. Such methods can produce internally consistent trajectories and ensemble spreads within the statistical manifold defined by the training data, but cannot properly generalize beyond it or respond to changes in physical boundary conditions or forcings that inject the trajectory out of the reference phase space. More research is needed to assess prediction skills and achievable forecast range of the method alone and in combination with high-resolution forecasting systems.
@@
The output of the global 1/4 (ORCA025-N4001) and 1/12 (ORCA0083-N006) resolution NEMO model used in this study is available on JASMIN ( https://jasmin.ac.uk/users/access/ ) from the following locations: /gws/nopw/j04/nemo_vol1/ORCA025-N401 and
/gws/nopw/j04/nemo_vol1/ORCA0083-N006 . Simulation data sets produced and analysed in this study as well as the software used in the analyses can be found in [26, 25].
@@ Acknowledgments
The author thanks the Natural Environment Research Council for the support of this work through the project ATLANTIS (P11742), as well as Andrew Coward and Chris Wilson for the production of and help with the NEMO datasets, respectively. The author thanks unknown referees for valuable comments and suggestions, which helped improve the manuscript